# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ C.3 Proof of Theorem 2: contraction of the population level EM iterates
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim d(\bar{\pi},\pi^{*})\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{\pi_{\ell}+\pi_{\ell}^{*}}{\bm{\pi}_{\ell}(\omega)}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\bm{\mu}_{\ell}(\omega))\right]\right\|_{\Sigma^{*}}$
$\lesssim d(\bar{\pi},\pi^{*})\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\bm{\mu}_{\ell}(\omega))\right]\right\|_{\Sigma^{*}}.$ (C.46)
@@
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\bm{\mu}_{\ell}(\omega))\right]\right\|_{\Sigma^{*}}$
$\leq\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N\right]\right\|_{2}+\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\sum_{b=1}^{L}\pi_{b}^{*}\mathbb{E}\left[I_{a\ell}(X,\omega)\left(\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right)\mid Y=b\right]\right\|_{\Sigma^{*}}$
$\leq\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\max_{b\in[L]}\left\|\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right\|_{\Sigma^{*}}.$
Using the fact that
[…]
we conclude
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\bm{\mu}_{\ell}(\omega))\right]\right\|_{\Sigma^{*}}\leq\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\left(\sqrt{\Delta_{\max}}+D_{\mu_{\ell}}\right).$ (C.48)
@@
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\left(\sqrt{\Delta_{\max}}+D_{\mu_{\ell}}\right)\right]~d(\bar{\pi},\pi^{*}).$ (C.49)
@@
$\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{k=1}^{L}{1\over\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[(X-\bm{\mu}_{\ell}(\omega))\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{k}},\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]\right\|_{\Sigma^{*}}$
$\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}{2\pi_{\ell}\over\bm{\pi}_{\ell}(\omega)}\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)(X-\bm{\mu}_{\ell}(\omega))\right]({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{\Sigma^{*}}$
@@
$~+\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}6\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)\sum_{b=1}^{L}1\{Y=b\}\right](\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega))({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{\Sigma^{*}}.$
@@
$\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}6\mathcal{R}_{I}^{(0)}\max_{b\in[L]}\left\|\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right\|_{\Sigma^{*}}\max_{k\in[L]}\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\|\bar{\mu}_{k}-\mu_{k}^{*}\|_{\Sigma^{*}}$
$\quad\lesssim\left(\sqrt{\Delta_{\max}}+D_{\mu_{\ell}}\right)\sqrt{\Delta_{\max}}~\mathcal{R}_{I}^{(0)}~d(\bar{M},M^{*})$
using (C.32), (C.47) and (C.44). We thus have
$\left\|\nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}~\lesssim~\left[\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\left(\sqrt{\Delta_{\max}}+D_{\mu_{\ell}}\right)\right]\sqrt{\Delta_{\max}}~d(\bar{M},M^{*}).$ (C.51)
@@
$\left\|\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{1}{\bm{\pi}_{\ell}(\omega)}\left\|\mathbb{E}\left[\left(X-\bm{\mu}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big{|}_{\omega},\bar{J}-J^{*}\right\rangle\right]\right\|_{\Sigma^{*}}.$
@@
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[(X-\bm{\mu}_{\ell}(\omega))\left\langle I_{a\ell}(X,\omega)\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)({\bm{e}}_{a}-{\bm{e}}_{\ell})^{\top},\bar{J}-J^{*}\right\rangle\right]\right\|_{\Sigma^{*}}$
$=\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\Sigma^{*-1/2}(X-\bm{\mu}_{\ell}(\omega))\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]\right\|_{2}.$
@@
$\Sigma^{*-1/2}(X-\bm{\mu}_{\ell}(\omega))\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}\stackrel{{\scriptstyle d}}{{=}}~NN^{\top}+T_{1}+T_{2}+T_{3}$ (C.52)
@@
$T_{3}$ $=\Sigma^{*-1/2}\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right)\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}.$
@@
$\leq\left|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N^{\top}\right]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\max_{b\in[L]}\left\|\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right\|_{\Sigma^{*}}$
$\leq 2\max_{b\in[L]}\left\|\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right\|_{\Sigma^{*}}\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}~d(\bar{J},J^{*})$
$\leq 2(D_{\mu_{\ell}}+\sqrt{\Delta_{\max}})\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}~d(\bar{J},J^{*})$
In the last step, we use (C.43) and (C.47). Similarly, by (C.39) and the argument in (C.3.2), we also find
@@
$=\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\max_{b\in[L]}\left\|\mu_{b}^{*}-\bm{\mu}_{\ell}(\omega)\right\|_{\Sigma^{*}}\left\|\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right\|_{\Sigma^{*}}2d(\bar{J},J)$
$\leq 4\sqrt{\Delta_{\max}}(D_{\mu_{\ell}}+\sqrt{\Delta_{\max}})\mathcal{R}_{I}^{(0)}d(\bar{J},J).$
@@
$\left\|\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim\left[\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}+(D_{\mu_{\ell}}+\sqrt{\Delta_{\max}})\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\sqrt{\Delta_{\max}}\mathcal{R}_{I}^{(0)}\right)\right]d(\bar{J},J^{*}).$ (C.53)
In view of Eqs. C.49, C.51 and C.53 and using $\mathcal{R}_{I}^{(1)}\leq\mathcal{R}_{I}^{(2)}$, we conclude that
$\begin{split}\left\|\bm{\mu}_{\ell}(\bar{\omega})-\bm{\mu}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}&\lesssim\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)~d(\bar{\omega},\omega^{*})+\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*})\\ &\qquad+\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*})+\mathcal{R}_{I}^{(0)}~d(\bar{\omega},\omega^{*})\right)D_{\mu_{\ell}}.\end{split}$ (C.54)
Note that the left hand side equals to $D_{\mu_{\ell}}$ by taking the supremum over $\bar{\omega}\in\mathcal{B}_{d}(\omega^{*})$. Since invoking Lemma 23 gives
[…]
with $\kappa_{w}$ given by (C.16), we obtain
$D_{\mu_{\ell}}$ $\leq\sup_{\bar{\omega}\in\mathcal{B}_{d}(\omega^{*})}2\left\{\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)d(\bar{\omega},\omega^{*})+\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*})\right\}\hskip-56.9055pt$
$\lesssim~\kappa_{w}\sqrt{\Delta_{\max}}\sup_{\bar{\omega}\in\mathcal{B}_{d}(\omega^{*})}d(\bar{\omega},\omega^{*})$
[…]
Plugging this into (C.54) gives
$\left\|\bm{\mu}_{\ell}(\bar{\omega})-\bm{\mu}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}\lesssim\kappa_{w}\sqrt{\Delta_{\max}}~d(\bar{\omega},\omega^{*}),$
thereby completing the proof of (C.18). ∎
@@ C.4 Key technical lemmas used in the proof of Theorem 20
Invoking (C.64) and (C.28) gives the result of $j=1$. The second term is bounded by
$\sup_{v\in\mathbb{S}^{d}}\mathbb{E}\left[v^{\top}NN^{\top}v1\{\mathcal{E}_{N}^{c}\}\right]$ $\leq\sup_{v\in\mathbb{S}^{d}}\sqrt{\mathbb{E}\left[(N^{\top}v)^{4}\right]}\sqrt{\mathbb{P}(\mathcal{E}_{N}^{c}\}}\leq 2\exp\Bigl{(}-\frac{c_{2}^{2}}{16c_{1}^{2}}\Delta_{\min}\Bigr{)}.$

### v2

@@ C.3 Proof of Theorem 2: contraction of the population level EM iterates
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim d(\bar{\pi},\pi^{*})\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{\pi_{\ell}+\pi_{\ell}^{*}}{\bm{\pi}_{\ell}(\omega)}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]\right\|_{\Sigma^{*}}$
$\lesssim d(\bar{\pi},\pi^{*})\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]\right\|_{\Sigma^{*}}.$ (C.45)
@@
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]\right\|_{\Sigma^{*}}$
$\leq\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N\right]\right\|_{2}+\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\sum_{b=1}^{L}\pi_{b}^{*}\mathbb{E}\left[I_{a\ell}(X,\omega)\left(\mu_{b}^{*}-\mu_{\ell}^{*}\right)\mid Y=b\right]\right\|_{\Sigma^{*}}$
$\leq\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}$
$\leq\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}.$ (C.46)
@@
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)~d(\bar{\pi},\pi^{*}).$ (C.47)
@@
$\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{k=1}^{L}{1\over\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[(X-\mu_{\ell}^{*})\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{k}},\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]\right\|_{\Sigma^{*}}$
$\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}{2\pi_{\ell}\over\bm{\pi}_{\ell}(\omega)}\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{\Sigma^{*}}$
@@
$~+\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}6\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)\sum_{b=1}^{L}1\{Y=b\}\right](\mu_{b}^{*}-\mu_{\ell}^{*})({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{\Sigma^{*}}.$
@@
$6\mathcal{R}_{I}^{(0)}\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}\max_{k\in[L]}\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\|\bar{\mu}_{k}-\mu_{k}^{*}\|_{\Sigma^{*}}\lesssim\Delta_{\max}~\mathcal{R}_{I}^{(0)}~d(\bar{M},M^{*})$
using (C.32) and (C.43). We thus have
$\left\|\nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}~\lesssim~\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)\sqrt{\Delta_{\max}}~d(\bar{M},M^{*}).$ (C.49)
@@
$\left\|\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{1}{\bm{\pi}_{\ell}(\omega)}\left\|\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big|_{\omega},\bar{J}-J^{*}\right\rangle\right]\right\|_{\Sigma^{*}}.$
@@
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[(X-\mu_{\ell}^{*})\left\langle I_{a\ell}(X,\omega)\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)({\bm{e}}_{a}-{\bm{e}}_{\ell})^{\top},\bar{J}-J^{*}\right\rangle\right]\right\|_{\Sigma^{*}}$
$=\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\Sigma^{*-1/2}(X-\mu_{\ell}^{*})\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]\right\|_{2}.$
@@
$\Sigma^{*-1/2}(X-\mu_{\ell}^{*})\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}\stackrel{{\scriptstyle d}}{{=}}~NN^{\top}+T_{1}+T_{2}+T_{3}$ (C.50)
@@
$T_{3}$ $=\Sigma^{*-1/2}\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\mu_{\ell}^{*}\right)\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}.$
@@
$\leq\left|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N^{\top}\right]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}$
Similarly, by (C.39) and the argument in (C.3.2), we also find
@@
$=\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}\left\|\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right\|_{\Sigma^{*}}2d(\bar{J},J)$
$\leq 4\Delta_{\max}\mathcal{R}_{I}^{(0)}d(\bar{J},J).$
@@
$\left\|\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim\left[\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}+\sqrt{\Delta_{\max}}\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\sqrt{\Delta_{\max}}\mathcal{R}_{I}^{(0)}\right)\right]d(\bar{J},J^{*}).$ (C.51)
In view of Eqs. C.47, C.49 and C.51 and using $\mathcal{R}_{I}^{(1)}\leq\mathcal{R}_{I}^{(2)}$, we conclude that
$\begin{split}\left\|\bm{\mu}_{\ell}(\bar{\omega})-\bm{\mu}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}&\lesssim\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)~d(\bar{\omega},\omega^{*})+\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*})\\ &\lesssim\kappa_{w}\sqrt{\Delta_{\max}}~d(\bar{\omega},\omega^{*})\end{split}$ (C.52)
where the last step invokes Lemma 22. This completes the proof of (C.18). ∎
@@ C.4 Key technical lemmas used in the proof of Theorem 19
Invoking (C.62) and (C.28) gives the result of $j=1$.